# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ 4.4 BP 17
In Brazil, the mechanism by which the public administration (the Union, states, municipalities, and autarchies) pays debts resulting from final judicial sentences is called the precatórios system (court order payment). This is a privileged payment system designed to protect the public treasures, regulated by Article 100 of the Constitution [^14] https://www.planalto.gov.br/ccivil_03/constituicao/Constituicao.htm#art100.. More precisely, precatórios are payment requisitions issued by the Judiciary in favor of individuals or legal entities that have won lawsuits against the public administration. These debts can be of an alimentary nature (such as salaries, pensions, retirements, indemnities for death or disability) or non-alimentary (other debts, such as expropriations and contracts).
@@
Article 100 establishes a set of rules on how precatórios should be paid. The first one concerns the order of payments: precatórios must be paid in the chronological order of presentation, respecting the division between alimentary and non-alimentary debts. There is a specific queue for each indebted public entity, meaning each entity has its own list of precatórios. However, alimentary precatórios have priority over non-alimentary ones. Elderly individuals (over 60 years old), those with serious illnesses, and people with disabilities also have priority in receiving amounts, up to a stipulated limit.
@@
Another important aspect of this system relates to the functioning of public budgets in Brazil. Public entities must forecast the necessary amounts for the payment of precatórios in their annual budgets, under penalty of federal, state, or municipal intervention, as the case may be. Thus, Paragraph 5 of Article 100 requires public entities to include in their annual budgets the necessary amounts for the payment of judicial precatórios presented by April 2 of each year [^15] This provision was originally located in Paragraph 1 of article 100 and the limit date for presenting precatórios was July 1st. These changes were introduced by Constitutional Amendments 62 and 114, as is further discussed ahead. These precatórios must be paid by the end of the following fiscal year. This rule ensures that debts arising from final judicial sentences are properly budgeted for and settled within a reasonable timeframe, preserving the real value owed to the creditor.
@@
Both issues — that of the order of payments and that of the processing of payments through the budget — become relevant because the precatórios system faces several challenges, such as delays in payments, especially in states and municipalities with a large volume of debts, and the constant need for constitutional amendments and adaptations to address default. The Binding Precedent 17 was created in November 2009 to clarify in which circumstances the Brazilian Public Treasuries needed to pay late payment interests (juros de mora). Specifically, the issue was whether this type of interest should apply to the period designated in the current Paragraph 5 (Paragraph 1, at the time of the BP’s creation) of article 100 of the Constitution. It was debated whether late payment interest would apply to the period between the deadline for presenting the precatórios and the actual date in which the payment is finally concluded by the Public Treasury. The BP states that, in this period, late payments are not applicable.
@@
The legal interpretation adopted by STF in BP 17 was based on the fact that the Public Treasuries were forced to process the payments through the budget. Thus, according to the Court, the time between the provision of the payment and the actual payment could not be considered a “delay” and should not be subjected to late payment interests.
@@
We draw the reader’s attention to the fact that only one month after the creation of the BP, Constitutional Amendment 62161616https://www.planalto.gov.br/ccivil_03/constituicao/Emendas/Emc/emc62.htm#art1 included a Paragraph 12 to Article 100, which establishes how late payment interests should be calculated, regardless of the period. Consequently, certain Justices understand this paragraph as overcoming the BP completely [^17] https://www.youtube.com/watch?v=7Rig_iyM8mY. However, it does not seem to be the prevalent understanding of the Court, as expressed by Justice Rosa Weber in the recurso extraordinário (extraordinary appeal) RE 577.465/2016. Many Justices still hold that late payment interests referred to in Paragraph 12 are only applicable if the payment does not occur by the end of the next fiscal year, subsequent to the provision. Then there would be an actual “delay” in the payment, not attributable to normal processing of payments through the budget.
Mathematical results
Fig. 10 presents the predictions of the five models for BP 17, together with a regex search for precatório and juros de mora (court orders, late payment interest). Some common trends stand out: a peak between 1995-1998 and between 2002-2005, as well as an increasing tendency starting from 2014, predicted by almost all the models. Besides, in addition to 1995-1998, regex detects two other peaks: in 2007 and 2015.
Models’ predictions for BP 17 Refer to caption
Figure 10: Number of documents predicted by each model for BP 17 in Dataset #2, represented as a histogram (window length of six months) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and #2 (solid), as described in LABEL:tab:adjusted_probabilities. Two views are given, the second one zooming in on the ordinate axis.
The difference between these predictions is directly explained by the words correlation in Fig. 11: except TF-IDF+randomforest, none of the models accurately captured the correlation of precatório (court order) with the others words (mora, late payment). In particular, the most important peak, between 2002-2005, is principally composed of documents mentioning “court order” but not “late payment”. This explains the absence of regex’s predictions in this period. This peak, even though not referring to cases directly linked to BP 17, contains relevant information, as studied below.
Another interesting observation is the significant correlation, in most of the models, between the terms moratório and juros/mora. The term moratório (moratorium) can, indeed, be used as a synonym for mora (late). However, it seems that the models have imposed that predicted documents when they contain the latter, must contain the former. This would explain why regex’s peak in 2007, mainly composed of documents citing juros de mora but not moratório, has not been detected by the models.
As with all the other binding precedents, BP 17 had random forest for the best model, in the sense that it predicts better the groundtruth curve, after the publication of the precedent. An additional search of words, given in Fig. 12, shows that during 2005-2009, almost no documents predicted by TF-IDF+randomforest satisfied the initial regex search. Indeed, in these documents, instead of the expression juros de mora are used the equivalent words juros and moratório. These are documents181818for instance, AI 525.183 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=620350 and AI 546.938-7 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=501102 for which BP 17 directly applies, highlighting the limitations of a naive regex search.
Additional metadata for documents predicted by TF-IDF+randomforest for BP 17 Refer to caption
Figure 12: Presence, in the documents predicted by the model TF-IDF+randomforest for BP 17, of the words precatório (court order), moratório (moratorium), juros (interests) and juros de mora (late payment interests).
Juridical discussion
The analysis above shows that the history of BP 17’s similar documents should be studied from both the models’ predictions — that may use synonyms instead of exact words from the precedent — and the regex-detected documents — that do not suffer from the correlations learned by the models. We note that they both share a small peak in 1993 (see Fig. 10), representing when the matter started to be taken to the court191919with a highlight to RE 149.466, in which Justice Sepúlveda Pertence concluded that juros de mora envolvem inadimplência, i.e., “late payment interests entail inability (to meet a financial obligation)”..
Starting with the models’ prediction, we give in Fig. 13 some additional information regarding TF-IDF+SVM. As it stands out, the 2002-2005 peak is almost exclusively composed of cases with origin in the state of São Paulo and of type Intervenção Federal (“Federal Intervention”). This is an exceptional measure of interference by the Union in the States, temporarily suppressing the autonomy of these entities. This event corresponds to the well-known refusal of São Paulo to pay its debts to alimentary/food creditors [^20] https://www.sandovalfilho.com.br/ha-20-anos-pedido-de-intervencao-federal-no-estado-de-sao-paulo-comecou-a-mudar-a-historia-dos-precatorios-judiciais/, prioritizing non-alimentary debts instead, in the early 2000s. A significant change occurred after a pivotal Supreme Court trial deciding on a Federal Intervention request in 2002, led by Attorney Antônio Roberto Sandoval Filho, and brought to trial by Justice Marco Aurélio. Though at the time there were questions about the claims brought before the court in this case — which, in fact, would not be successful —, this decision marked a turning point, drawing attention to the unsustainable nature of the court-ordered debt situation in Brazil. Even though we still observe a high value on the curve from 2016 (Figure 13), the payment of debts in São Paulo was considered maximum priority by the Court of Justice of the State of São Paulo [^21] https://www.tjsp.jus.br/Imprensa/Campanhas/Precatorios, with more than 19 billion reais (4 billion dollars) being released for court-ordered payments in 2023222222https://www.tjsp.jus.br/Noticias/Noticia?codigoNoticia=95996&pagina=1.
Additional metadata for documents predicted by TF-IDF+SVM for BP 17 Refer to caption
@@
Figure 13: First: Type of legal process used in the documents predicted by the model TF-IDF+SVM for BP 17, among RE (Recurso Extraordinário), ARE (Recurso Extraordinário com Agravo) and IF (Intervenção Federal). Second: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Third: Presence of the words precatório (court order), moratório (moratorium) and juros (interests).
We now turn to regex’s predictions, with Fig. 14 representing additional metadata. We notice a 2006-2007 peak regarding militares (“military personnel”) in the State of Rio de Janeiro. A search shows that this peak mainly consists of documents employing the word militares which may be traced to some issues with wages [^23] https://www.conjur.com.br/2004-jan-28/militares_rio_nao_receber_reajuste_1307/. This peak seems to coincide with a problem that Rio de Janeiro was dealing with in 2004, regarding the adjustment of military wage. At that time, civil public servants received a wage adjustment to overcome inflation, and Justice decided that military public servants would not be included. Further analyses are necessary to assess whether this is the case for the predicted documents, however, our methodology is useful in identifying these patterns and justifying these analyses.
Additional metadata for documents detected by regex for BP 17 Refer to caption
Figure 14: Top: State to which the cases satisfying the regex query belong, with the label “UNKNOWN” when the metadata is unavailable. Bottom: Presence of the words militares (militaries) and 4.357.
@@
In 2015, the peak of documents satisfying the regex query can be explained by ADIs 4.357242424https://portal.stf.jus.br/processos/detalhe.asp?incidente=3813700 and 4.425252525https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=5067184 (Ação Direta de Inconstitucionalidade, Direct Action for Unconstitutionality). This relates to a very specific, technical legal discussion, regarding a transition rule introduced by Constitutional Amendment 62. The Court found Constitutional Amendment 62 to be partially unconstitutional, due to some of the rules it introduced for the order of payments — specially some of the provisions concerning preference due to old age — and, as a result, there was uncertainty regarding how the interest rate for late payments should be calculated during the transition period.
@@
The data suggests that BP 17 was partially ineffective in reducing the number of cases brought to STF. The BP was created after the major peek of cases concerning its main subject had already subsided. Soon after its creation, major changes to the underlying constitutional text were introduced by Constitutional Amendments 62 and 114, raising questions if the BP was still standing. Furthermore, by the end of the period of analysis, we may see again a tendency towards an increase in the number of cases concerning the subject of the BP. The only peek identified by the regex query after creation of the BP, in 2015, can be attributed to an external event — matters related to the effects of the declarations of unconstitutionality in ADIs 4.357 and 4.425. This corresponds to our hypothesis External conjecture not directly related to the precedent.
@@
But the main lesson that can be drawn from this case seems to be that the dynamic nature of the legal system sometimes will bring changes that render a BP ineffective, or that raise questions about its validity soon after it is issued. It might be argued that, in a Civil Law system, the preferred solution to a problem of legal interpretation or implementation comes from a change in the underlying legal texts — and not through Case Law or a BP. Nevertheless, it is frequently necessary to rely on judicial solutions as they provide a faster solution to urgent problems and address concrete problems more directly. In the case of BP 17, however, we see how the dynamics of the legal system, how changes of the legal text and disputes over related topics can limit the effectiveness of this judicial tool to address legal problems and reduce litigation.
@@ 4.5 BP 26
In December 2009, Binding Precedent 26 was introduced to address two key legal issues, regarding individuals convicted of heinous crimes (crimes hediondos e equiparados), a category that encompasses offenses like torture, homicide, drug dealing, or rape. As specified in the 1990’s Law n°8.072262626Law 8.072https://www.planalto.gov.br/ccivil_03/leis/l8072.htm, a more severe treatment was imposed on these prisoners, in particular prohibiting progress through the three prison regimes, closed, semi-open, and open (Article 2, Paragraph 1). In HC 82.959/SP [^27] HC 82.959/SP https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=ac&docID=79206&pgI=156&pgF=160, the court found this provision to be unconstitutional. This sparked two subsequent discussions related to the determination of regime progression that would be answered in BP 26. Firstly, BP 26 clarified whether the law 11.464/07282828Law 11.464 https://www.planalto.gov.br/ccivil_03/_Ato2007-2010/2007/Lei/L11464.htm, allowing sentence progression for heinous crimes, applied retroactively. Secondly, it determined whether judges could require a criminological examination (exame criminológico) for sentence progression. The following wording was established, which would address both issues raised:
@@
It is worth mentioning that the debate that led to the voting of BP 26 reveals a high level of criticism among public defenders. The dispute, especially led by the Public Defender’s Office of the State of São Paulo, represented by Dr. Rafael Ramia Muneratti, focused on the second part of the binding precedent, regarding the criminological examination. It was argued that this examination, extensively rejected by psychologists and other specialists [^29] The contradiction between the conduct of criminological examination and the Psychologists’ Code of Ethics is discussed in https://www.crpsp.org/noticia/view/1846/nota-tecnica-sobre-a-suspensao-da-resolucao-cfp-0122011---atuacao-dao-psicologao-no-ambito-do-sistema-prisional., would contradict the princípio da individualização da pena (principle of individualization of punishment). The next year, this disagreement was explicitly written by the Public Defender’s Office under the name “Thesis 53” [^30] Tese 53 da defensoria Pública do Estado de São Paulo https://www2.defensoria.sp.def.br/dpesp/Conteudos/Materia/MateriaMostra.aspx?idItem=61257&idModulo=9706..
Mathematical results
We show in Fig. 15 the predictions of the five models, as well as a regex search for the words exame criminológico and progressão de regime (criminological examination, regime progression). Except for TF-IDF+randomforest, they all present a peak in 2006, as well as an increasing tendency starting from 2010. Although exhibiting similar behaviors, the five models differ a lot regarding the number of documents estimated: for 2018, LSTM estimates approximately five times as many documents as TF-IDF+SVM or TF-IDF+logistic.
Models’ predictions for BP 26 Refer to caption
Figure 15: Number of documents predicted by each model for BP 26 in Dataset #2, represented as a histogram (window length of six months) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and #2 (solid), as described in LABEL:tab:adjusted_probabilities. Two views are given, the second one zooming in on the ordinate axis.
Moreover, a manual inspection shows that the documents estimated by LSTM after 2010 are of poor quality, in the sense that they do not fit in the context of BP 26. To explain this discrepancy, we illustrate in Fig. 16 the correlations between the top TF-IDF words obtained in LABEL:tab:importance_features: criminológico, execuções, observará, progressão, realizacão and regime (criminological, executions, will observe, progression, realization, regime). It is worth noting the significant correlation between criminológico and progressão in the groundtruth image. All the models show a similar correlation. To complement this data, we give in Fig. 17 the frequency of citations of the words. As observed, LSTM tends to predict many documents solely based on the word regime. In fact, among the 324,756 documents of Dataset #2 that employ this word, the model predicts 5328 documents. In comparison, TF-IDF+SVM only predicts 1574. We could say that LSTM tends to detect positive cases that present the word regime, taking little account of its conjunction with the context of BP 26. This inevitably results in a very large number of documents, given the importance of the theme of regime progression for prisoners.
Correlations between words among the models’ predictions for BP 26 Refer to caption
Figure 16: Correlations between the selected relevant words, when restricted to documents citing BP 26 (“Groundtruth”), or documents predicted by one of our models.
Frequency of words among the models’ predictions for BP 26 Refer to caption
Figure 17: Frequency of the selected relevant words, when restricted to documents citing BP 26 (“Groundtruth”), or documents predicted by one of our models.
Another interesting observation is the quantity of documents detected by all the models, except TF-IDF+randomforest, during the peak between 2006 and 2008. Similarly, the regex search finds 659 documents in this period. As it turns out, a manual inspection of these documents shows that they are indeed related to the theme of BP 26 and, hence are considered correct predictions of the models. A further inspection shows that, among the most important features of the random forest (see LABEL:tab:importance_features), the words realizacão (realization) and subjetivos (subjective) are present in respectively 83.7% and 80.7% of the 241 predicted documents. In opposition, they are only present in 60.8% and 60.4% of the 659 documents detected by regex between 2006 and 2008. This suggests that this TF-IDF model has learned unnecessary restrictions, imposing the presence of words realizacão and subjetivos, resulting in too few predicted documents. In what follows, we will only consider the documents given by the models TF-IDF+SVM and TF-IDF+logistic, as these seem to be the most reliable estimations.
Juridical discussion

### v3

@@ 5.4 BP 17
In Brazil, the mechanism by which the public administration (the Union, states, municipalities, and autarchies) pays debts resulting from final judicial sentences is called the precatórios system (court order payment). This is a privileged payment system designed to protect the public treasures, regulated by Article 100 of the Constitution [^20] Article 100 https://www.planalto.gov.br/ccivil_03/constituicao/Constituicao.htm#art100.. More precisely, precatórios are payment requisitions issued by the Judiciary in favor of individuals or legal entities that have won lawsuits against the public administration. These debts can be of an alimentary nature (such as salaries, pensions, retirements, indemnities for death or disability) or non-alimentary (other debts, such as expropriations and contracts).
@@
Article 100 establishes rules for the payment of precatórios. Payments must follow the chronological order of presentation, with a distinction between alimentary and non-alimentary debts. Each indebted public entity maintains its own list of precatórios. Alimentary precatórios take precedence over non-alimentary ones. Additionally, elderly individuals (over 60 years old), those with serious illnesses, and people with disabilities are given priority, up to a stipulated limit.
@@
Another important aspect of this system relates to the functioning of public budgets in Brazil. Public entities must forecast the necessary amounts for the payment of precatórios in their annual budgets, under penalty of federal, state, or municipal intervention, as the case may be. Thus, Paragraph 5 of Article 100 requires public entities to include in their annual budgets the necessary amounts for the payment of judicial precatórios presented by April 2 of each year [^21] This provision was originally located in Paragraph 1 of Article 100 and the limit date for presenting precatórios was July 1st. These changes were introduced by Constitutional Amendments 62 and 114, as is further discussed ahead.. These precatórios must be paid by the end of the following fiscal year. This rule ensures that debts arising from final judicial sentences are properly budgeted for and settled within a reasonable timeframe, preserving the real value owed to the creditor.
@@
Both issues — that of the order of payments and that of the processing of payments through the budget — become relevant because the precatórios system faces several challenges, such as delays in payments, especially in states and municipalities with a large volume of debts, and the constant need for constitutional amendments and adaptations to address default. The Binding Precedent 17 was created in November 2009 to clarify in which circumstances the Brazilian Public Treasuries need to pay late payment interests (juros de mora). Specifically, the issue was whether this type of interest should apply to the period designated in the current Paragraph 5 (Paragraph 1, at the time of the BP’s creation) of Article 100 of the Constitution. It was debated whether late payment interest would apply to the period between the deadline for presenting the precatórios and the actual date on which the payment is finally concluded by the Public Treasury. The BP states that, in this period, late payments are not applicable.
@@
The legal interpretation adopted by the STF in BP 17 was based on the fact that the Public Treasuries were forced to process the payments through the budget. Thus, according to the Court, the time between the provision of the payment and the actual payment could not be considered a “delay” and should not be subjected to late payment interests.
@@
We draw the reader’s attention to the fact that only one month after the creation of the BP, Constitutional Amendment 62 included a Paragraph 12222222Constitutional Amendment 62 (see Art.1º/§12) https://www.planalto.gov.br/ccivil_03/constituicao/Emendas/Emc/emc62.htm#art1 to Article 100, which establishes how late payment interests should be calculated, regardless of the period. Consequently, certain Justices understand this paragraph as overcoming the BP completely [^23] Overview of the controversies surrounding BP 17 https://www.youtube.com/watch?v=7Rig_iyM8mY. However, it does not seem to be the prevalent understanding of the Court, as expressed by Justice Rosa Weber in the Recurso Extraordinário (Extraordinary Appeal) RE 577.465/2016242424RE 577.465 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=11795802. Many Justices still hold that late payment interests referred to in Paragraph 12 are only applicable if the payment does not occur by the end of the next fiscal year, after the provision, for there would be an actual “delay” in the payment, not attributable to normal processing of payments through the budget.
Discussion
The prediction curves studied in Section 4.2 (see Fig. 12) show that the history of BP 17’s similar documents should be studied from both the models’ predictions — that may use synonyms instead of exact words from the precedent — and the regex-detected documents — that do not suffer from the correlations learned by the models.
Starting with the models’ prediction, we give in Fig. 19 some additional information regarding Longformer. In chronological order, the first peak is visible between 1995-1997. Manual inspection indicates some abundance of complaints by public entities mostly in cases of land expropriation. For, after the promulgation of the newest Brazilian Constitution, in 1988, it was established by Article 33252525Act of Transitory Constitutional Dispositions (see Art. 33) https://www2.camara.leg.br/legin/fed/conadc/1988/constituicao.adct-1988-5-outubro-1988-322234-normaatualizada-pl.pdf of the Act of Transitory Constitutional Dispositions (Ato das Disposições Constitucionais Transitórias) that, except for alimentary, precatórios could be paid by equal installments through eight years. Because installments were said to be equal, no delay interests were predicted. Judges from a lower court of appeals in the State of São Paulo, however, decided that compensations due to land expropriations were not contemplated by the text of Article 33 and, therefore, late interests should be applicable. This generated a series of appeals to the STF, where judges decided that the exception to the non-applicability of Article 33 was restricted to alimentary precatórios, which did not include cases of expropriation (see the search for “art. 33” at the bottom of the figure). Article 33’s effect naturally died as the period of eight years predicted in its text ended. It is noteworthy that a less significant effect from the same period, but that was also detected by both Longformer, regex, and TF-IDF+SVM and logistic regression, corresponds to appeals, again mostly by public entities, related to the applicability of late payment interests, this time to alimentary precatórios. However, we could not identify a single event causing this effect, although the appeals come, in their majority, from the States of São Paulo and Paraná.
Additional metadata for documents predicted by Longformer for BP 17 Refer to caption
@@ Discussion
Figure 19: Top: Type of legal process used in the documents predicted by the model Longformer for BP 17, among AI (Agravo de Instrumento), RE (Recurso Extraordinário), and IF (Intervenção Federal). Middle: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Bottom: Presence of the words precatório (court order payment), juros de mora (late payment interests), 9.494/97 (the law), and art. 33 (the article). The bins are six months long.
We now turn to the main peak, between 2002-2005. As it stands out, it is almost exclusively composed of cases with origin in the State of São Paulo and of type Intervenção Federal (Federal Intervention). This is an exceptional measure of interference by the Union in the States, temporarily suppressing the autonomy of these entities. This event corresponds to the well-known refusal of São Paulo to pay its debts to alimentary/food creditors [^26] Report on the request for Federal Intervention in SP https://www.sandovalfilho.com.br/ha-20-anos-pedido-de-intervencao-federal-no-estado-de-sao-paulo-comecou-a-mudar-a-historia-dos-precatorios-judiciais/, prioritizing non-alimentary debts instead, in the early 2000s. A significant change occurred after a pivotal Supreme Court trial deciding on a Federal Intervention request in 2002, led by Attorney Antônio Roberto Sandoval Filho, and brought to trial by Justice Marco Aurélio. Though at the time there were questions about the claims brought before the court in this case — which, in fact, would not be successful —, this decision marked a turning point, drawing attention to the unsustainable nature of the court-ordered debt situation in Brazil. Even though we still observe a high value on the curve from 2016 (Fig. 19), the payment of debts in São Paulo was considered maximum priority by the Court of Justice of the State of São Paulo [^27] TJSP’s precatórios campaign https://www.tjsp.jus.br/Imprensa/Campanhas/Precatorios, with more than 19 billion reais (4 billion dollars) being released for court-ordered payments in 2023282828TJSP’s disclosure https://www.tjsp.jus.br/Noticias/Noticia?codigoNoticia=95996&pagina=1.
Regarding the last peak, in 2007-2008, the analyzed documents predominantly address legal disputes involving military personnel and the application of late payment interest in cases of delayed compensation, often resulting in precatórios. Central to these cases is Article 1º-F of Law 9.494/1997292929Law 9.494/1997, Article 1º-F https://www.planalto.gov.br/ccivil_03/leis/l9494.htm, which establishes a reduced annual interest rate of 6% for public debts, instead of 12% (see “9.494/97” at the bottom of Fig. 19). More precisely, the article has been introduced by Provisional Measure 2.180-35303030MPV 2.180-35 https://www.planalto.gov.br/ccivil_03/LEIS/L9494.htm#art1f. in 2001 (later rectified in 2009). Military personnel challenged this reduced rate, arguing it violates the constitutional principle of isonomy (equality before the law), particularly when compared to civilian public servants or private-sector norms. While many rulings uphold the lower rate as consistent with fiscal policy, others question its fairness, especially given the unique service conditions of military personnel. These cases highlight a systemic conflict between fiscal governance, the equitable treatment of public servants, and the adequacy of precatório-related compensation mechanisms.
@@
We now turn to regex’s predictions, with Fig. 20 representing additional metadata. The peaks in 1995-1997 and 2007-2008 have already been identified in Longformer’s predictions (also visualized in Fig. 20), so we skip ahead to 2015. Between 2014 and 2017, the number of documents satisfying the regex query derive from ADIs 4.357313131ADI 4.357 https://portal.stf.jus.br/processos/detalhe.asp?incidente=3813700 and 4.425323232ADI 4.425 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=TP&docID=5067184 (Ação Direta de Inconstitucionalidade, Direct Action for Unconstitutionality). This relates to a very specific, technical legal discussion, regarding a transition rule introduced by Constitutional Amendment 62 (already mentioned). The Court found this amendment to be partially unconstitutional, due to some of the rules it introduced for the order of payments — especially some of the provisions concerning preference due to old age — and, as a result, there was uncertainty regarding how the interest rate for late payments should be calculated during the transition period.
Additional metadata for documents predicted by regex for BP 17 Refer to caption
Figure 20: Top: State to which the cases satisfying the regex query belong, with the label “UNKNOWN” when the metadata is unavailable. Bottom: Presence of the words art. 33 (Article 33), militares (military personnel) and 4.357 (the law). The bins have a length of six months.
@@
The data suggests that BP 17 was partially ineffective in reducing the number of cases brought to the STF. The BP was created after the major peak of cases concerning its main subject had already subsided. Soon after its creation, major changes to the underlying constitutional text were introduced by Constitutional Amendments 62 and 114, raising questions if the BP was still standing. Furthermore, by the end of the period of analysis, we may see again a tendency towards an increase in the number of cases concerning the subject of the BP. The only peak identified by regex after the creation of the BP, in 2015, can be attributed to an external event — matters related to the effects of the declarations of unconstitutionality in ADIs 4.357 and 4.425. This corresponds to our hypothesis External conjecture not directly related to the precedent.
@@
But the main lesson that can be drawn from this case seems to be that the dynamic nature of the legal system sometimes will bring changes that render a BP ineffective, or that raise questions about its validity soon after it is issued. It might be argued that, in a Civil Law system, the preferred solution to a problem of legal interpretation or implementation comes from a change in the underlying legal texts — and not through Case Law or a BP. Nevertheless, it is frequently necessary to rely on judicial solutions as they provide a faster solution to urgent problems and address concrete problems more directly. In the case of BP 17, however, we see how the dynamics of the legal system, and how changes in the legal text and disputes over related topics can limit the effectiveness of this judicial tool to address legal problems and reduce litigation.
@@ 5.5 BP 26
In December 2009, Binding Precedent 26 was introduced to address two key legal issues, regarding individuals convicted of heinous crimes (crimes hediondos e equiparados), a category that encompasses offenses like torture, homicide, drug dealing, and rape. As specified in the 1990’s Law 8.072333333Law 8.072 https://www.planalto.gov.br/ccivil_03/leis/l8072.htm, more severe treatment was imposed on these prisoners, in particular prohibiting progress through the three prison regimes, closed, semi-open, and open (Article 2, Paragraph 1). In HC 82.959/SP [^34] HC 82.959/SP https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=ac&docID=79206&pgI=156&pgF=160, the court found this provision to be unconstitutional. This sparked two subsequent discussions related to the determination of regime progression that would be answered in BP 26. Firstly, BP 26 clarified whether the law 11.464/07353535Law 11.464 https://www.planalto.gov.br/ccivil_03/_Ato2007-2010/2007/Lei/L11464.htm, allowing sentence progression for heinous crimes, applied retroactively. Secondly, it determined whether judges could require a criminological examination (exame criminológico) for sentence progression. The following wording was established, which would address both issues raised:
@@
It is worth mentioning that the debate that led to the voting of BP 26 reveals a high level of criticism among public defenders. The dispute, especially led by the Public Defender’s Office of the State of São Paulo, represented by Dr. Rafael Ramia Muneratti, focused on the second part of the binding precedent, regarding the criminological examination. It was argued that this examination, extensively rejected by psychologists and other specialists [^36] The contradiction between the conduct of criminological examination and the Psychologists’ Code of Ethics is discussed in https://www.crpsp.org/noticia/view/1846/nota-tecnica-sobre-a-suspensao-da-resolucao-cfp-0122011---atuacao-dao-psicologao-no-ambito-do-sistema-prisional., would contradict the princípio da individualização da pena (principle of individualization of punishment). The following year, this disagreement was explicitly formalized by the Public Defender’s Office under the name “Thesis 53” [^37] Thesis 53 of the Public Defender’s Office of the State of São Paulo https://www2.defensoria.sp.def.br/dpesp/Conteudos/Materia/MateriaMostra.aspx?idItem=61257&idModulo=9706..
Discussion